OpenStudy (anonymous):

what is the area of the circle with circumference of 9.42 inches

OpenStudy (anonymous):

@Heatherlynnlee You need to work backwards.

OpenStudy (anonymous):

You need to find the radius, but that can be found by using the Circumferance formula.

OpenStudy (anonymous):

C=2πr

OpenStudy (anonymous):

My bad, C = 2(pi)r

OpenStudy (anonymous):

So we're going to divide by pi(2), which is 6.28

OpenStudy (anonymous):

9.42 / 6.28 = 1.48

OpenStudy (anonymous):

Now we need to do 1.48 / 2 . Then we should get our radius.

OpenStudy (alexandervonhumboldt2):

7.065

OpenStudy (alexandervonhumboldt2):

9.42/3.14/2=1.5 1.5*1.5*3.14=7.065

OpenStudy (anonymous):

Yep. Was my explaining correct?

OpenStudy (anonymous):

It's a bit tricky as well, because it doesn't specify whether rounding is ~ tenth or ~ hundreth.